OpenStudy (anonymous):

You sit down to take a true-or-false test with 7 questions. If you randomly guess on all questions, how many possible outcomes are there for the 7-question test?

ganeshie8 (ganeshie8):

this is similar to tossing 7 coins

ganeshie8 (ganeshie8):

you can get either a `Head` or a `Tail` on each coin - `2` possibilities for each coin

ganeshie8 (ganeshie8):

Since you have a total of 7 coins, total number of outcomes = \(2\times 2 \cdots (\text{7 times }) = 2^7\)

OpenStudy (anonymous):

you have 2 choices per question and 7 questions for test 2 ^ 7 = 128

OpenStudy (mathstudent55):